OpenStudy (anonymous):

What is the difference of f and g? f(x) = –20x g(x) = 2x + 9

OpenStudy (anonymous):

Just subtract them: -20x - 2x + 9

OpenStudy (anonymous):

(f – g)(x) = –22x + 9 (f – g)(x) = –22x – 9 (f – g)(x) = –18x + 9 (f – g)(x) = 22x + 9

OpenStudy (anonymous):

Just subtract -20x - 2x

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

-22x + 9?

Nnesha (nnesha):

difference subtract so  first distribute